OpenStudy (anonymous):

The pair of points are inversely related. (9, a) and (6, 12) MEDAL AND FAN REWARD

OpenStudy (anonymous):

help @KinzaN

OpenStudy (anonymous):

I'm not sure how to work through this. If they're inversely related then (a, 9) is the point on the graph where (6, 12) is also a point.

OpenStudy (whpalmer4):

if you have an inverse relationship, y = k/x for some value of k. we can find that value of k from (6,12): 12 = k/6 solve for k 6*12 = k/6*6 72 = k Now, we know that y = 72/x, and we know that x = 9. plug in x = 9 and the value of y will be your answer for a.